OpenStudy (anonymous):

A stuntwoman (m=60kg) jumps off the roof of a building, and then safely lands on a pile of cardboard boxes, which takes 0.4s to crumple and bring her to a complete stop. The force of the cardboard boxes on the stuntwoman is 2000N and the area of the force vs. time graph that was provided to me was 600 N/s. What was the stuntwoman's momentum and velocity at t=0 seconds (just prior to hitting the boxes)? HELP ME

OpenStudy (ujjwal):

\[Force=\frac{\Delta P}{\Delta t}\]where \(\Delta P\)= change in momentum \(\Delta t\) is time of impact Let P be initial momentum. Since the body finally comes to rest. change in momentum=P so, here, 2000 = P/0.4 or, P=800 Also, P=mv therefore, v=P/m=800/60=13.33 m/s

OpenStudy (anonymous):

thx :)